Let R = {(1,4), (2,1), (2,5),(2,4),(4,3),(5,3),(3,2)} on the set A = {1, 2, 3, 4, 5}. Use Warshall’s algorithm to find transitive closure of R.


1
Expert's answer
2021-10-29T02:50:01-0400

Let "A=\{1,2,3,4,5\}" and "R" be the relation on "A" such that "R= \{(1,4),(2,1),(2,5),(2,4),(4,3),(5,3),(3,2)\}." 


Warshall's Algorithm:


Step 1: put "W:=M_R,\ k:=0."


Step 2: put "k:=k+1."


Step 3: for all "i\ne k" such that "w_{ik}=1," and for all "j" put "w_{ij}:=w_{ij}\lor w_{kj}."


Step 4: if "k=n" then STOP else go to the step 2.


Let us find the transitive closure of "R" by Warshall's Algorithm:


"W^{(0)}=M_R=\begin{pmatrix}
0 & 0 & 0 & 1 & 0\\
1 & 0 & 0 & 1 & 1\\
0 & 1 & 0 & 0 & 0\\
0 & 0 & 1 & 0 & 0\\
0 & 0 & 1 & 0 & 0
\end{pmatrix}"



"W^{(1)}=\begin{pmatrix}
0 & 0 & 0 & 1 & 0\\
1 & 0 & 0 & 1 & 1\\
0 & 1 & 0 & 0 & 0\\
0 & 0 & 1 & 0 & 0\\
0 & 0 & 1 & 0 & 0
\end{pmatrix}"


"W^{(2)}=\begin{pmatrix}
0 & 0 & 0 & 1 & 0\\
1 & 0 & 0 & 1 & 1\\
1 & 1 & 0 & 1 & 1\\
0 & 0 & 1 & 0 & 0\\
0 & 0 & 1 & 0 & 0
\end{pmatrix}"


"W^{(3)}=\begin{pmatrix}
0 & 0 & 0 & 1 & 0\\
1 & 0 & 0 & 1 & 1\\
1 & 1 & 0 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1
\end{pmatrix}"


"W^{(4)}=\begin{pmatrix}
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1
\end{pmatrix}"


"M_{R^*}=W^{(5)}=\begin{pmatrix}
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1\\
1 & 1 & 1 & 1 & 1
\end{pmatrix}"


It follows that the transitive closure "R^*" of "R" is the following:


"R^*=A\times A."